OpenStudy (anonymous):

simplify sqrt 50 x^5 y^3 z^6. show work.

OpenStudy (anonymous):

\[ \sqrt{50 x^5 y^3 z^6} = 5x^2yz^3\sqrt{2xy} \]

OpenStudy (anonymous):

Assuming x,y,z are positive reals.

OpenStudy (anonymous):

can u draw it

OpenStudy (saifoo.khan):

Online colllege?

OpenStudy (anonymous):

And they don't allow you to use latex? That sucks

OpenStudy (anonymous):

online algebra 2

OpenStudy (saifoo.khan):

EASY as cake!

OpenStudy (anonymous):

1okay then answer it :) thanks

OpenStudy (anonymous):

lol how can this be college? thats like 7th grade middle school algebra

OpenStudy (saifoo.khan):

that guy just answerd.

OpenStudy (saifoo.khan):

lol, UoP is just like that.

OpenStudy (anonymous):

what is UoP?

OpenStudy (anonymous):

university of phoenix???

OpenStudy (saifoo.khan):

yea.

OpenStudy (anonymous):

god gracious..

OpenStudy (anonymous):

i know the answer it there but im looking for the work.. no need to be rude. u guys r annoying bye.

OpenStudy (saifoo.khan):

srsly? sorry for that please dont leave.

OpenStudy (saifoo.khan):

we can help you if you want steps!!

OpenStudy (anonymous):

oh sorry :-( wasnt meaning to be rude